# Partonic Entropy of the Proton from DGLAP Evolution

### v1

@@ Abstract
We investigate the concept of partonic entropy of the proton within the Dokshitzer–Gribov–Lipatov–Altarelli–Parisi (DGLAP) evolution scheme of collinear parton distributions. We show that such entropy increases monotonically with the evolution scale. The mechanism underlying this growth is illustrated using a simplified model of DGLAP evolution, which highlights the importance of including saturation effects at small $x$ in the evolution of parton distributions. Based on existing literature, we present two simplified models of parton saturation at small $x$. In one of these models, partonic entropy is identified with entanglement entropy and proposed as an observable to be tested experimentally.
@@ I Introduction
The rapid growth of gluon distribution function for the Bjorken variable $x\to 0$, found in the deep inelastic scattering (DIS) experiments at the HERA accelerator, stimulated enormous theoretical activity. This behavior had been anticipated by perturbative quantum chromodynamics (pQCD) prior to the first data from HERA, through the all-order resummation of large logarithms $\log(1/x)$. These developments led to computational frameworks based on evolution equations such as the DGLAP [1, 2, 3] or BFKL (Balitsky-Fadin-Kuraev-Lipatov) equations [4, 5]. The resulting physical picture describes the nucleon as a dense system of partons (gluons and sea quark–antiquark pairs) probed in the small-$x$ regime of DIS. A central feature of this regime is parton saturation, by which the rapid growth of parton densities at small-$x$ is moderated by gluon recombination effects. This phenomenon induces non-linear corrections to evolution equations that are linear in parton distribution functions (PDFs) [6].
In recent years, there has been a growing interest in understanding the entropy of parton systems within the nucleon, as investigated through DIS [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Of particular interest is the dense system of partons with small momentum fractions $x$, where saturation occurs. Several interesting proposals have been made in this context, the most intriguing of which is a conjecture relating the entropy of the produced hadrons to the entanglement entropy of the parton system probed in DIS at small $x$ [10]. This means that the ideas known from the studies of the foundation of quantum mechanics are successfully applied to the high-energy particle scattering. Moreover, quantum information tools are employed to characterize dense partonic systems [24, 26].
@@
The paper is organized as follows. In Section II, we define partonic entropy and prove that it is positive. In Section III, we present the DGLAP evolution equations in the form most appropriate for proving that the partonic entropy is a monotonic function of the evolution parameter, which is done in Section IV. In Section V, we relate the partonic entropy to the relative entropy known from the theory of information. In Section VI, we present a simplified model of DGLAP evolution to illustrate that the saturation of the small-$x$ behavior of PDFs is essential for taming the unbounded growth of the partonic entropy with the evolution scale. For completeness, in Section VII, we present the dipole model developed in the literature to realize parton saturation, while in Section VIII, we discuss a similar model in which the concept of entanglement entropy in DIS was introduced.
@@ II Partonic entropy
No minus sign is introduced on the r.h.s. because $S(t)$ is non-negative, in contrast to the entropy density,
@@
which is negative for $P(x,t)<1$. The non-negativity of $S(t)$ can be readily proven using a well-known inequality
@@
and due to the normalization condition (3), one obtains
@@
Note that the non-negativity is also a consequence of the integration range $x\in[0,1]$. For different integration domains, the generally defined entropy (5) may admit a negative lower bound. This is in contrast to relative entropy, to be discussed later, which is intrinsically non-negative.
@@ VI Entropy growth interpretation
This observation raises a question about saturation effects which involve not only gluon splitting but also gluon recombination [34, 35, 36, 37, 38, 39, 40]. The latter phenomenon suppresses the small-$x$ increase of the parton distributions and hopefully imposes an upper bound on the partonic entropy, see the recent discussion in [41] in the context of the Color Glass Condensate (CGC) [6], an effective theory of gluon saturation.
@@
Therefore, the partonic entropy $S$ is no longer singular in the considered limit and has an upper bound which saturates its vale. Notice hat the small-$x$ contribution (48) also saturates the momentum sum rule since in the limit $\alpha\to 0^{+}$
@@ VII Reaction-diffusion model of parton saturation
For the completeness of the discussion, we will present simplified models of parton saturation effects at small $x$, discussed in the literature, which illustrate the saturation of entropy as $x\to 0$. Strictly speaking, this is a discussion of the Regge limit of QCD, in which the hard scale $Q^{2}$ is kept fixed while the invariant energy of the partonic processes $s\to\infty$. Thus, the relevant quantities are parton distributions that depend not only on parton’s momentum fraction $x$ but also on its transverse momentum $k_{\perp}$. However, in the simplified models presented below, the latter dependence is neglected.
@@ VIII Entanglement entropy in DIS
$\ln\,(xG(x))=S_{\text{h}adron}$ (72)
This identification has been tested against experimental data in [14, 16, 21, 23] with a positive result. An application of a generalized version of the model (61) to the description of hadron multiplicities in the hadron-hadron scattering data was presented in [28]. A precise formulation of the entanglement idea is still lacking, although the studies in this direction were carried out in [20, 19, 18] in two-dimensional QCD.
@@ Acknowledgements.
We thank Krzysztof Kutak and Anna Staśto for useful discussions. This work was supported by the National Science Center, Poland, Grants Nos. 2019/33/B/ST2/02588 and 2024/53/B/ST2/00968.

### v2

@@ Abstract
We investigate the concept of partonic entropy of the proton within the Dokshitzer–Gribov–Lipatov–Altarelli–Parisi (DGLAP) evolution scheme of collinear parton distributions. We show that such entropy increases monotonically with the evolution scale. The mechanism underlying this growth is illustrated using a simplified model of DGLAP evolution, highlighting the importance of including saturation effects at small $x$ in the evolution of parton distributions, which tame the otherwise unbounded growth of partonic entropy. Based on existing literature, we present two simplified models of parton saturation at small $x$. In one of these models, partonic entropy is identified with entanglement entropy and proposed as an experimentally testable observable.
@@ I Introduction
The rapid growth of gluon distribution function for the Bjorken variable $x\to 0$, found in the deep inelastic scattering (DIS) experiments at the HERA accelerator, stimulated enormous theoretical activity. This behavior had been anticipated by perturbative quantum chromodynamics (pQCD) prior to the first data from HERA, through the all-order resummation of large logarithms $\log(1/x)$. These developments led to computational frameworks based on evolution equations such as the DGLAP [1, 2, 3] or BFKL (Balitsky-Fadin-Kuraev-Lipatov) equations [4, 5]. The resulting physical picture describes the nucleon as a dense system of partons (gluons and sea quark–antiquark pairs) probed in the small-$x$ regime of DIS. A central feature of this regime is parton saturation, by which the rapid growth of parton densities at small-$x$ is moderated by gluon recombination effects. This phenomenon induces non-linear corrections to evolution equations that are linear in parton distribution functions (PDFs). The effective QCD-based theory of parton saturation, known as the Color Glass Condensate, is presented in a comprehensive review [6].
In recent years, there has been a growing interest in understanding the entropy of parton systems within the nucleon, as investigated through DIS or hadron-hadron scattering [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Of particular interest is the dense system of partons with small momentum fractions $x$, where saturation occurs. Several interesting proposals have been made in this context, the most intriguing of which is a conjecture relating the entropy of the produced hadrons to the entanglement entropy of the parton system probed in DIS at small $x$ [10]. This means that the ideas known from the studies of the foundation of quantum mechanics are successfully applied to the high-energy particle scattering [30]. Moreover, quantum information tools are employed to characterize dense partonic systems [24, 26].
@@
The paper is organized as follows. In Section II, we define partonic entropy and prove that it is positive. In Section III, we present the DGLAP evolution equations in the form most appropriate for proving that the partonic entropy is a monotonic function of the evolution parameter, which is done in Section IV. In Section V, we relate the partonic entropy to the relative entropy known from the theory of information. In Section VI, we present a simplified model of DGLAP evolution to illustrate that the saturation of the small-$x$ behavior of PDFs is essential for taming the unbounded growth of the partonic entropy with the evolution scale. For completeness, in Section VII, we present the dipole model developed in the literature to realize parton saturation, while in Section VIII, we discuss a similar model in which the concept of entanglement entropy in high energy scattering was introduced.
@@ II Partonic entropy
In Eq. (4), the minus sign is necessary to ensure positivity of $S$, since $0<P_{n}\leq 1$ and the logarithm is always negative. This is not the case for the probability distribution $P(x,t)$, which has no the upper bound; consequently the logarithm in Eq. (5) can be either positive or negative. In particular, when $P(x,t)<1$, the entropy density,
@@
is negative. However, no minus sign is introduced on the right-hand side of Eq. (5). because $S(t)$ is always non-negative.
This can be readily proven using a well-known inequality
@@
and due to normalization condition (3), one obtains
@@
Note that the non-negativity is also a consequence of the integration range $x\in[0,1]$. For different integration domains, the generally defined entropy (5) may admit a negative lower bound. This is in contrast to relative entropy, to be discussed later, which is always non-negative.
@@ VI Entropy growth interpretation
This observation raises a question about saturation effects which involve not only gluon splitting but also gluon recombination [37, 38, 39, 40, 41, 42, 43]. The latter phenomenon suppresses the small-$x$ increase of the parton distributions and hopefully imposes an upper bound on the partonic entropy, see [7] and the recent discussion in [44] in the context of the Color Glass Condensate (CGC) [6], an effective theory of gluon saturation.
@@
Therefore, the partonic entropy $S$ is no longer singular in the considered limit and has an upper bound which saturates its value. Notice hat the small-$x$ contribution (48) also saturates the momentum sum rule since in the limit $\alpha\to 0^{+}$
@@ VII Reaction-diffusion model of parton saturation
For the completeness of the discussion, we will present simplified models of parton saturation effects at small $x$, discussed in the literature, which illustrate the saturation of entropy as $x\to 0$. Strictly speaking, this is a discussion of the Regge limit of QCD, in which the hard scale $Q^{2}$ is kept fixed while the invariant energy of the partonic processes $\sqrt{s}\to\infty$. Thus, the relevant quantities are parton distributions that depend not only on parton’s momentum fraction $x$ but also on its transverse momentum $k_{\perp}$. However, in the simplified models presented below, the latter dependence is neglected.
@@ VIII Entanglement entropy in DIS
$\ln\,(xg(x))\simeq S_{\text{h}adron}$ (72)
This identification has been improved by including quark distributions and tested against experimental data in [14, 16, 21, 23] with a positive result. An application of a generalized version of the model (61) to the description of hadron multiplicities in the hadron-hadron scattering data was presented in [28].
A precise formulation of the entanglement idea is still lacking, although the studies in this direction were carried out in [20, 19, 18] in two-dimensional QCD; see also [31] for a recent, comprehensive review of the entanglement in high energy scattering.
@@ Acknowledgements.
We thank Krzysztof Kutak, Stanisław Mrówczyński and Anna Staśto for useful discussions. This work was supported by the National Science Center, Poland, Grants Nos. 2019/33/B/ST2/02588 and 2024/53/B/ST2/00968.